(cot 2 a (sec a – 1))/1 + sin a = sec 2 a( 1 – sin a)/(1 + sec a)

(cot2 a (sec a – 1))/1 + sin a = sec2 a( 1 – sin a)/(1 + sec a)

Grade:12th pass

1 Answers

Akshay
8 years ago
write cot2a = cos2a/sin2a = (1-sin2a)/(1-cos2a) = (1+sin a)(1-sin a) / (1+cos a)(1-cos a).
Put cos a = 1/sec a. cot2a = (1+sin a)(1- sin a)*sec2a/(sec a+1)*(sec a – 1).
Put this in LHS and you will get the result